three equal point charges Q are arranged at the corners of a square of side L. draw the situation. what is the potential at the fourth corner, taking V=0 at a great distance?
1
Expert's answer
2019-08-21T10:34:52-0400


The electric potential generated by a point charge is given by:


"V=K\frac{Q}{r}"


Where:

  • Coulomb constant "K"
  • Electric charge "Q"
  • Distance from the load to the vertex."r"

All charges have the same magnitude but the distance varies.


  • Distance of electric charge 1 to point P.


"r_{1}=L"


  • The electric potential generated by charge 1 at point P is:


"V_{1}=K\frac{Q}{r_{1}}"


Now


"V_{1}=K\frac{Q}{L}"


  • Distance of electric charge 2 to point P


The Pythagorean Theorem is applied.


"r_{2}=\sqrt{L^{2}+L^{2}}"


"r_{2}=\sqrt{2L^{2}}"


FInally "r_{2}=\sqrt{2}L"


Electric potential generated from charge 2 to point P


"V_{2}=K\frac{Q}{r_{2}}"


Now


"V_{2}=K\frac{Q}{\sqrt{2}L}" rationalizing "V_{2}=K\frac{Q}{\sqrt{2}L}*\frac{\sqrt{2}}{\sqrt{2}}"


Finally "V_{2}=\sqrt{2}K\frac{Q}{2L}"


  • Distance from charge 3 to point P


"r_{3}=L"


Electric potential generated by charge 3 at point P


"V_{3}=K\frac{Q}{r_{3}}"


Now


"V_{3}=K=\frac{Q}{L}"


The potential at point P is


"V_{P}=V_{1}+V_{2}+V_{3}"


Now


"V_{P}=K\frac{Q}{L}+\sqrt{2}K\frac{Q}{2L}+K\frac{Q}{L}"


Adding similar terms


"V_{P}=2K\frac{Q}{L}+\sqrt{2}K\frac{Q}{2L}"


Adding fractions


"V_{p}=\frac{4KQ+\sqrt{2}KQ}{2L}"


Factoring.


"V_{p}=\frac{4+\sqrt{2}}{2}\frac{KQ}{L}"